OpenStudy (anonymous):

Okay, so the instructions say to use Standard Form to write an equation for the line described->y= 4/5x-1 that means I know the slope is 4/5, how do I find out an ordered pair?

OpenStudy (saifoo.khan):

ordered pair = (x,y)

OpenStudy (anonymous):

so it would be (4/5 , 1)

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

but the question dosnt mentions that for what thing u want the ordered pair.

OpenStudy (anonymous):

huh? well from what weve been learning we need an ordered pair--- another problem on the hw was slope=3, containing (4, -8)

OpenStudy (anonymous):

but on the problem im on, it just gave me y=4/5x-1

OpenStudy (saifoo.khan):

Look, u just wrote an ordered pair, (4, -8)

OpenStudy (anonymous):

i know.. ur point being... lol.. point.. get it!;P

OpenStudy (saifoo.khan):

if u want any ordered pair, then there can be infinte answrs. y = 4x/5 - 1 Set x = 0 y = 4(0)/5 - 1 y = -1 So ordered pair = (x, y) = (0, -1)

OpenStudy (karatechopper):

use intercepts! from tht equation u can get one (0,-1) and then find te x int

OpenStudy (anonymous):

BUT it makes no since!

OpenStudy (anonymous):

sense

OpenStudy (saifoo.khan):

LOl, then u do a thing, listen to those "freaking nerds".

OpenStudy (anonymous):

oh yeah... you know what, im just gonna go ask my dad real quick

OpenStudy (anonymous):

Ah ah, nonno i said frikkin, not freaking, frikkin!

OpenStudy (karatechopper):

wait a min u dont need an ordered pair for this! i can write this in standard from!

OpenStudy (anonymous):

-4/5x+y=-1? yeah, uh no we have to SHOW our work... ugh bleh

OpenStudy (karatechopper):

i will show the work

OpenStudy (anonymous):

gee thanks, im really lost:P

OpenStudy (karatechopper):

and tht is not standard form: in standard form there cant be decimals or fractions and x cant be negative

OpenStudy (karatechopper):

so u got started right! but ended wrong, lets start from where u ended which is: -4/5x+y=-1

OpenStudy (anonymous):

yeah

OpenStudy (karatechopper):

k so lets start

OpenStudy (anonymous):

doesnt it have to be negative tho, since u subtracted.. i mean i know it cant be for standard form, but where did the - sign go?

OpenStudy (karatechopper):

-4/5x+y=-1 lets make x positive first: so if we make x positive the others have to be treated as well 4/5x-y=1 now lets make there be no fractions or decimals (4/5x-y=1)5 distributive property 4/5x(5) - y(5)= 1(5) 4x-5y=5: is ur new standard form equation

OpenStudy (karatechopper):

i explained where the neg would go.. if u dont understand tht i can tell u y with an example

OpenStudy (anonymous):

okay, where did u get the 5 in the multipl. part?

OpenStudy (karatechopper):

well if u have a frac and u want to get rid of it, then u want to multiply the GCF

OpenStudy (karatechopper):

er srry meant LCM

OpenStudy (karatechopper):

and in this case it was 5

OpenStudy (anonymous):

i seeeeeeeeee thanks soo muchhh(:

OpenStudy (karatechopper):

welcome got anymore questions?

OpenStudy (anonymous):

um i think im good

OpenStudy (karatechopper):

k and if u wanna check if u r right just ask

OpenStudy (anonymous):

cool man cool